OpenStudy (anonymous):

HELP! please/thanks :) Jason wants to make an 11% alcohol solution. he has already poured 9 gallons of pure water into a beaker. how many gallons of 20% alcohol must he add to this to create the desired mixture?

OpenStudy (0bucchr):

solve with simultaneous equations: y is total solution x is added alcohol at 20% 0.2*x + 9 = 0.11*y y = 9+ x